Circular No. 6364

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1996 B2 (HYAKUTAKE) A. Dutrey, Institut de Radio Astronomie Millimétrique (IRAM); D.

Despois, Observatoire de Bordeaux; D. Bockelée-Morvan, P. Colom, N. Biver, J. Crovisier, E. Gérard, and H. Rauer, Observatoire de Paris; M. Grewing, S. Guilloteau, R. Lucas, R. Neri and J. Wink, IRAM, report the detection of CH_{3}CN in comet C/1996 B2 on Mar. 26.4 UT at the IRAM Plateau de Bure Interferometer using its four antennas in a single-dish mode: “The CH_{3}CN 5(1)-4(1) 91.9853-GHz and 5(0)-4(0) 91.9871-GHz lines were detected with integrated line area on a TA* scale of 0.028 ± 0.004 and 0.025 ±0.004 K km/s, respectively. Assuming a rotation temperature of 80 K, we derive a CH_{3}CN-production rate of 1.8 × 10^{25} mol/s. For a water-production rate of 3.5 × 10^{29} mol/s (IAUC 6355), the inferred CH_{3}CN abundance is 5 × 10^{−5}.”

H. Böhnhardt, Institute for Astronomy and Astrophysics, University of Munich, reports on CCD observations by O. Bärnbandner and himself made during Mar. 23.90–24.18 UT with the 0.80-m telescope (+ broadband B, V, and R filters) at the Mt. Wendelstein Observatory: “After image analysis using adaptive Laplace filtering, we detected time-variable jets close to the nucleus and shell-like structures in the sunward coma hemisphere. The shell structures are apparently identical with the remnants of old jets from previous revolutions of the nucleus. We can identify up to three generations of signatures from on-going and previous jet activities in our 3′.6 × 4′.5 field of view: the actual (0) level with jets close to the nucleus (up to 2–4 jets at a time, typical extension of some 10′′–40′′ from the nucleus; curvature mostly counterclockwise, occasionally also straight, but confined to the sunward coma hemisphere); another shell level ∼ 1′ away from the nucleus in the coma hemisphere; and a third, weaker shell ∼ 2′ away from the center. The features described are most likely produced by dust. The projected shell expansion velocity was estimated to be of the order of 250 m/s. Examples of the coma jets and shells can be found under

http://www.eso.org/educnpubrelns/phot-hya-mar25-hb.html.”

Further selected naked-eye m_{1}, coma-diameter, and tail-length estimates: Mar. 27.10 UT, 0.4, ∼ 60′, 45^{o} (H. Mikuž, Ljubljana, Slovenia); 27.43, 1.0, 1^{o}.0, 70^{o} (A. Hale, Cloudcroft, NM); 28.15, 0.7, –, > 50^{o} (M. Tichý, Kleť, Czech Republic); 28.90, 0.5, 45′ (J.-C. Merlin, Le Creusot, France); 29.18, 1.5, 40′, 10^{o} (C. E. Spratt, Victoria, BC; strong moonlight); 29.81, 1.3, 40′, 20^{o} (K. Hornoch, Lelekovice, Czech Republic).

1996 March 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green